OpenStudy (wade123):

CALCULUS HELP PLEASE (:

OpenStudy (wade123):

sorry to be bothering everyone with these questions, ive just kept these to the side because people havnt been responding on OS and i just dont get how to solve these problems, i appreciate the help(:

OpenStudy (wade123):

@ybarrap please help !

OpenStudy (ybarrap):

Does this make sense? $$ a(t)={dv(t)\over dt}\\ \int a(t)dt=\int dv(t)\\ v(t)=\int a(t)dt=\int (2t+\sin t) dt $$ Can you do this?

OpenStudy (wade123):

yes! im going to do it now, could you check me after?

OpenStudy (wade123):

i got v(t)=t^2+cos(t)+3 @ybarrap

OpenStudy (wade123):

@perl

OpenStudy (anonymous):

Check your integration of the sine function. \[ \int sin(t) dt = -cos(t) + C\] ^_^

OpenStudy (ybarrap):

Close, $$ v(t)=\int a(t)dt=\int (2t+\sin t) dt=t^2-\cos t + C $$ Since you know that \(v(0)=4\) you can find C: $$ v(0)=4=0^2-\cos 0 + C\\ \implies C=4+1=5 $$ So now you have everything! Make sense?